# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v2

@@ 3.2 Data generation of microscopic distributions
We generate synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy. Each simulation has a time span of $250\,\text{s}$. To reduce computational storage costs while retaining essential temporal features, we subsample the resulting pedestrian trajectories at intervals of $\delta t=0.25\,\text{s}$. As already discussed, we generate pedestrian trajectories from various initial spatial distributions modelling diverse crowd scenarios. In particular, we consider $C=20$ different initializations with zero velocities and positions distributed via: (i) uniform distributions to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) “double bell-shaped” Gaussian distributions to capture the formation of two distinct crowd groups and (iv) sinusoidal distributions to generate periodic macroscopic behavior. Details on the form and parameters of the above initializations are provided in Appendix B, distinguishing between those used for training and testing the proposed framework in Tables B.1 and B.2, respectively. The resulting datasets are comprised of 20 cases with different initializations, each case containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1000}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1000$ time steps correspond to the samples recorded from each trajectory. Thus, the total datasets over all cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(20\cdot 1000)}$ used for training and testing, respectively.
@@ 3.3 Extraction of continuous density fields
For the first step of our framework, we derive macroscopic density fields from pedestrian positions via KDE, as described in Section 2.2.1. To achieve this, we discretize the computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behavior, while minimizing artificial grid effects. For the KDE implementation on the cell centroids, we choose Gaussian kernels in Eq. (6) tuned with bandwidths $\mathbf{H}=diag(h_{x},h_{y})=(3,2)$ for suppressing noise and preserving features of the microscopic data, while ensuring that local structures remain clear without introducing spurious artifacts. To account for the physical obstacle in $\Omega$, a binary mask is applied to explicitly zero out density estimation on the cell centroids of the obstacle geometry. To account for the periodic boundary conditions, we further perform domain augmentation for the KDE support. All the details regarding KDE implementation are provided in Appendix C.
Using the above procedure, we derive the density fields $\rho^{(c)}\big(x_{i},y_{j},t_{k}\big)$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing $\mathcal{X}_{ts}$ data sets. For conforming to the assumptions of Proposition 1, we finally normalize each density field following Eq. (19), thus constructing the normalized density matrices $X_{tr},X_{ts}\in\mathbb{R}^{n_{c}\times n_{t}}$ in the form of Eq. (9), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=20\cdot 1000$.
@@ 3.4 POD for the restriction and lifting operators
For constructing the restriction and lifting operators in Eqs. (20) and (21) respectively, we employ POD on the training dataset $X_{tr}$ to determine the latent dimension $d$ and the POD basis $U_{d}$. Using the economy-size SVD (which reduces computational complexity in thin matrices while maintaining exact reconstruction capability), we determine $d$ by the best low-rank approximation of $X_{tr}$ retaining $99\%$ of the energy; thus, computing the minimum number $d$ satisfying the criterion:
@@
where $r=\min(n_{c},n_{t})=1600$ is the rank of $X_{tr}$, and $\sigma_{i}>0$ for $i=1,\ldots,r$ are the singular values in descending order. Next, using the first $d$ left singular vectors of $X_{tr}$, we form $U_{d}$ for completing the construction of the restriction and lifting operators in Eqs. (20, 21). To assess the accuracy of restriction and lifting operators, we evaluate the POD reconstruction error on the training and testing data sets by the relative $L_{2}$ error:
@@
where $x(t_{k})\in X_{tr},X_{ts}$ with $k=1,\ldots,K$ is the normalized density defined in Eq. (19).
@@ 3.5 ROMs for latent space dynamics and their training
Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct ROMs for learning the dynamics in the latent space. For this purpose, we learn linear MVAR and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}(t_{k-1})\}$. To conform to this sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq. (20) on the training data $X_{tr}$.
@@
As a preliminary step, we first tune the lag window $w$, within a bound of $w_{\text{max}}=100$, using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [93, 94, 95]; details are provided in Appendix D. Both criteria are likelihood-based, with AIC tending to select larger windows than BIC. We note that BIC and AIC are well-suited for MVARs models since they assume stationarity of the data and white noise residuals. For comparison, we adopt the same $w$ for learning LSTMs, as the one tuned for MVARs by these criteria.
@@
For the MVARs model training, we first verify that all input time series are stationary using the augmented Dickey–Fuller (ADF) test [96] with a threshold of significance level $p<0.01$. Then, we compute the coefficients $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{j}$ in Eq. (22) using ordinary least-squares (OLS) to solve the optimization problem in Eq. (24).
@@
For the LSTM model, we consider a single LSTM layer with $16$ units and a hyperbolic tangent activation function in the hidden layers; the output layer is linear, as denoted in Eq. (28). To train the LSTM model, we solve the optimization problem in Eq. (29) using the Adam optimizer [91] with automatic differentiation for the gradients of the loss function, enabled through MATLAB’s Deep Learning Toolbox [97]. The model’s parameters are initialized with a Glorot uniform initialization [98], the initial learning rate is set to $l_{r}=10^{-3}$, and the mini-batch size is set to $32$; the latter hyperparameters were tuned on a trial-and-error basis, providing a good compromise between computational efficiency and generalization without exceeding memory constraints. The maximum number of epochs was set to $200$ as no further improvement was achieved using more epochs. To further avoid overfitting of the LSTM, we split the training data into training and validation sets using a 80-20% ratio, and shuffle each data set without breaking the internal temporal structure of the sequences. We compute the loss function value on the validation set for every epoch and stop the training after 5 consecutive validation checks with no improvement.
@@
$MSE_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\rVert^{2},$ (34)
where $\hat{\boldsymbol{y}}_{c}(t_{k})$ is the prediction in Eq. (26), to enable direct comparison with the MSE-based loss function of LSTMs. Additionally, we report the computational times required for training both ROMs (for LSTMs, we report the mean error and 10–90% error percentiles statistics across the 50 runs); we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the OLS method employed for solving the related optimization problem.
@@ 3.6 Accuracy assessment of the proposed framework
Having constructed/trained each step of the proposed framework–construction of the restriction/lifting operators and training of the ROMs–, we finally assess the forecasting performance of the whole framework on the $C=20$ different initializations of the testing set $\mathcal{X}_{ts}$; see Appendix B for details. In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$, using Eq. (19). Then, following the proposed restrict-evolve with ROM-lift framework in Eq. (5), we forecast the normalized density at the $k$ time step, using the $w$ past normalized densities in both an open-loop and closed-loop setting. For the open-loop setting evaluation, we compute the prediction:
$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(x^{(c)}(t_{k-1})),\ldots,\mathcal{R}(x^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (35)
while for the closed-loop setting evaluation, we compute the prediction:
$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(\hat{x}^{(c)}(t_{k-1})),\ldots,\mathcal{R}(\hat{x}^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (36)
where $\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators in Eqs. (20) and (21) and $\Phi_{\text{ROM}}$ is the latent dynamics ROM with parameters $\mathbf{p}$ that can be either the trained MVAR or LSTM model. In particular, for the LSTM model, we consider the parameter set resulting in the best training performance among the 50 runs. Notice that the open-loop prediction in Eq. (35) considers the restricted values of the observed normalized densities $\mathcal{R}(x^{(c)}(t_{k-j}))$ for $j=1,\ldots,w$ as inputs, while the closed-loop prediction in Eq. (36) instead considers the predicted–from previous steps–normalized densities $\mathcal{R}(\hat{x}^{(c)}(t_{k-j}))$ in a recursive fashion. Since the open-loop setting requires the “ground truth” densities, we use it here to evaluate the one-step predictions of our framework. For the desired, in practice, longer predictions, we utilize the closed-loop setting, which enables us to recursively feed subsequent predictions into Eq. (36), resulting in an autoregressive time-stepping framework also considering the effect of one-step forecast error accumulation.
To quantify the end-to-end prediction accuracy, we compute the relative, reconstructed in the ambient-space, $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ (either for the open-loop or the closed-loop setting) per case $c$ in the testing set and time step $k$ as:
@@
For a summary of statistics, we report the mean and the 10–90% percentiles of the above errors over all cases and time steps. To track down the prediction accuracy in time, we further depict the mean and 10–90% percentiles of the error evolution in time over all cases.
All numerical computations were executed on a PC equipped with an 11th Gen Intel(R) Core(TM) i7-1165G7 2.80 GHz CPU and 16 GB of system memory.

### v3

@@ 3.2 Data generation of microscopic distributions
We generated synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy. Each simulation has a time span of $250\,\text{s}$. To reduce computational storage costs while retaining essential temporal features, we subsample the resulting pedestrian trajectories at intervals of $\delta t=0.25\,\text{s}$. As already discussed, we generate pedestrian trajectories from various initial spatial distributions modelling diverse crowd scenarios. In particular, we consider $C=20$ different initializations with zero velocities and positions distributed via: (i) uniform distributions to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) “double bell-shaped” Gaussian distributions to capture the formation of two distinct crowd groups and (iv) cosinusoidal distributions to generate periodic macroscopic behavior. Details on the form and parameters of the above initializations are provided in C, distinguishing between those used for training and testing the proposed framework in Tables C.1 and C.2, respectively.
For the unidirectional flow, the resulting datasets are comprised of 20 cases with different initializations, each containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1000}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1000$ time steps correspond to the samples recorded from each trajectory. Thus, the total datasets over all cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(20\cdot 1000)}$ used for training and testing, respectively.
For the counterflow, we generated data from the same set of $C=20$ initializations; see C. Here, the initialization for group 1 is identical to that of the unidirectional case (with half the agents), while group 2 is initialized symmetrically to group 1 along $x=24\,\text{m}$, preserving the $y$-positions. Each initialization case now produces two separate data matrices $\mathcal{X}^{(1,c)}$ for group 1 and $\mathcal{X}^{(2,c)}$ for group 2, both of dimension $\mathbb{R}^{100\times 1000}$. Consequently, the complete training and testing datasets for the counterflow are $\{\mathcal{X}_{tr}^{(1)},\mathcal{X}_{tr}^{(2)}\}$ and $\{\mathcal{X}_{ts}^{(1)},\mathcal{X}_{ts}^{(2)}\}$, respectively.
@@ 3.3 Extraction of continuous density fields
For the first step of our framework, we derive macroscopic density fields from pedestrian positions via KDE, as described in Section 2.2.1. To achieve this, we discretize the corridor computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behavior, while minimizing artificial grid effects. For the KDE implementation on the cell centroids, we choose Gaussian kernels in Eq. (6) tuned with bandwidths $\mathbf{H}=diag(h_{x},h_{y})=(3,2)$ for suppressing noise and preserving features of the microscopic data, while ensuring that local structures remain clear without introducing spurious artifacts. To account for the physical obstacle in $\Gamma\subset\Omega$, we applied a binary mask to the KDE-estimated density field, enforcing zero density in $\Gamma$. This is necessary because KDE can produce non-zero density estimates near obstacle boundaries when pedestrians are in close proximity. To account for the periodic boundary conditions, we further perform domain augmentation for the KDE support. All the details regarding KDE implementation are provided in B.
Using the above procedure, in the unidirectional flow case, we derived the density fields $\rho^{(c)}\big(x_{i},y_{j},t_{k}\big)$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing $\mathcal{X}_{ts}$ data sets. For conforming to the assumptions of Proposition 1, we normalize each density field following Eq. (33), thus constructing the normalized density matrices $X_{tr},X_{ts}\in\mathbb{R}^{n_{c}\times n_{t}}$ in the form of Eq. (9), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=20\cdot 1000$. In the counterflow case, we followed the same procedure to extract and normalize the density fields of each group separately, as it is required by Proposition 2, yielding the matrices $X_{tr}^{(1)},X_{ts}^{(1)}\in\mathbb{R}^{n_{c}\times n_{t}}$ for group 1 and $X_{tr}^{(2)},X_{ts}^{(2)}\in\mathbb{R}^{n_{c}\times n_{t}}$ for group 2.
@@ 3.4 POD for the restriction and lifting operators
We construct the restriction and lifting operators by employing POD on the training data, determining the latent dimension $d$ and the POD basis $U_{d}$ via the economy-size SVD. In all cases, we determine $d$ by the best low-rank approximation retaining $99\%$ of the energy; thus, computing the minimum number satisfying the criterion:
@@
where $r=\min(n_{c},n_{t})=1600$ is the rank of the data matrix, and $\sigma_{i}>0$ for $i=1,\ldots,r$ are the singular values in descending order. The basis $U_{d}$ is then formed by the first $d$ left singular vectors.
For the unidirectional flow case, we applied this procedure to the training dataset $X_{tr}$ and the resulting basis $U_{d}$ is then used to construct the restriction and lifting operators in Eqs. (34) and (35), respectively. For the counterflow case, the above procedure is done per group, to compute the separate POD bases $U_{d_{1}}^{(1)}$ and $U_{d_{2}}^{(2)}$ from the group-specific training datasets $X_{tr}^{(1)}$ and $X_{tr}^{(2)}$, respectively. Following Proposition 2, we then computed the SVD of the cross-covariance matrix $C^{(12)}=\frac{1}{n_{t}}\bar{X}^{(1)}_{tr}\left(\bar{X}_{tr}^{(2)}\right)^{\top}$ of the training sets, which is required to form the augmented bases $\tilde{U}_{d_{l}}^{(l)}$ (for $l=1,2$) defined in Eq. (20). These augmented bases were subsequently used to construct the group-specific restriction and lifting operators in Eqs. (36) and (37).
For both configurations, to assess the accuracy of restriction and lifting operators, we evaluate the POD reconstruction error on the training and testing data sets by the relative $L_{2}$ error:
@@
where $x(t_{k})\in X_{tr},X_{ts}$ denotes the normalized density in Eq. (33) at time $t_{k}$ for $k=1,\ldots,K$. For the counterflow, this error is computed separately for each group using the corresponding operators and data sets $X^{(l)}_{tr},X^{(l)}_{ts}$ for $l=1,2$.
@@ 3.5 ROMs for latent space dynamics and their training
Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct ROMs for learning the dynamics in the latent space. For this purpose, we learn linear MVAR and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$.Note that in the counterflow, a single ROM is trained on the concatenated latent vectors from both groups $\boldsymbol{y}_{c}(t_{k})=[\boldsymbol{y}^{(1)}_{c}(t_{k}),\boldsymbol{y}^{(2)}_{c}(t_{k})]^{\top}$, as detailed in Section 2.2.3. To conform to sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq. (34) on the training data $X_{tr}$ (for the unidirectional case) and via Eq. (36) on $X_{tr}^{(1)},X_{tr}^{(2)}$ (for the counterflow case).
@@
As a preliminary step, we first tune the lag window $w$, within a bound of $w_{\text{max}}=20$, using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [81, 2, 100]; details are provided in D. Both criteria are likelihood-based, with AIC tending to select larger windows than BIC. We note that BIC and AIC are well-suited for MVARs since they assume stationarity of the data and white noise residuals. For comparison, we adopt the same $w$ for learning LSTMs, as the one tuned for MVARs by these criteria.
@@
For the MVAR model training in both configurations, we first verified that all input time series are stationary using the augmented Dickey–Fuller (ADF) test [39] with a threshold of significance level $p<0.01$. Then, we computed the coefficients $\boldsymbol{A}_{j}$ in Eq. (38) using ordinary least squares to solve the optimization problem in Eq. (40). For the counterflow case, where the related data-matrix was ill-conditioned, we employed ridge regularization with $\lambda=10^{-6}$.
@@
For the LSTM model in both configurations, we considered a single LSTM layer with $16$ units and a hyperbolic tangent activation function in the hidden layers; the output layer is linear, as denoted in Eq. (44). To train the LSTM model, we solve the optimization problem in Eq. (45) using the Adam optimizer [74] with automatic differentiation for the gradients of the loss function, enabled through MATLAB’s Deep Learning Toolbox [107]. The model’s parameters are initialized with a Glorot uniform initialization [52], the initial learning rate is set to $l_{r}=10^{-3}$, and the mini-batch size is set to $32$; the latter hyperparameters were tuned on a trial-and-error basis, providing a good compromise between computational efficiency and generalization without exceeding memory constraints. The maximum number of training epochs was set to $40$ for the unidirectional flow case and to $70$ for the counterflow case, based on the saturation of the training loss, which yielded no additional improvement beyond these epochs.
@@
$MSE_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert^{2}_{2},$ (50)
where $\hat{\boldsymbol{y}}_{c}(t_{k})$ is the prediction in Eq. (42), to enable direct comparison with the MSE-based loss function of LSTMs. Additionally, we report the computational times required for training both ROMs (for LSTMs, we report the mean error and 10–90% error percentiles statistics across the 50 runs); we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the method employed for solving the related optimization problem.
@@ 3.6 Accuracy assessment of the proposed framework
Having constructed/trained each step of the proposed framework (construction of the restriction/lifting operators and training of the ROMs), we finally assess the forecasting performance of the whole framework on the $C=20$ different initializations of the testing set $\mathcal{X}_{ts}$; see C for details. In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$, using Eq. (33). While ROMs are trained in an open-loop setting using ground truth embeddings, we evaluate the proposed framework in a closed-loop setting (i.e., recursively feeding subsequent predictions to the autoregressive ROMs), as this reflects the intended use in practice. In particular, following the restrict-evolve with ROM-lift framework in Eq. (5), we compute the closed-loop prediction:
$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\hat{\boldsymbol{y}}^{(c)}(t_{k})\right),\qquad\hat{\boldsymbol{y}}^{(c)}(t_{k})=\Phi_{\text{ROM}}\left(\hat{\boldsymbol{y}}^{(c)}(t_{k-1}),\ldots,\hat{\boldsymbol{y}}^{(c)}(t_{k-w});\mathbf{p}\right),$ (51)
where $\hat{\boldsymbol{y}}^{(c)}(t_{k-j})$ are latent points predicted recursively by $\Phi_{\text{ROM}}$ in the previous $j=1,\ldots,w$ steps. Closed-loop forecasting requires the initial $j=1,\ldots,w$ embeddings of the observed densities $\mathcal{R}(x^{(c)}(t_{j}))$. Here,
$\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators (defined in Eqs. (34,35) for the unidirectional flow or in Eqs. (36, 37) for the counterflow), and $\Phi_{\text{ROM}}$ is the latent dynamics ROM with parameters $\mathbf{p}$ that can be either the trained MVAR or LSTM model; for LSTMs, we select the best-performing model among the 50 training runs. For completeness, we provide the open-loop setting and corresponding one-step prediction results in E and F for the unidirectional flow and counterflow cases, respectively. In open-loop, the latent points are obtained from ground truth embeddings $\boldsymbol{y}^{(c)}(t_{k-j})=\mathcal{R}(x^{(c)}(t_{k-j}))$ rather than previous predictions $\hat{\boldsymbol{y}}^{(c)}(t_{k-j})$ in Eq. (51).
For evaluation purposes, the prediction is obtained at every time step $k$. However, in practice, the lifting is applied on-demand only when a high-dimensional reconstruction is required. This renders the scheme particularly efficient, as the evolution is governed by the ROM in the latent space.
To quantify the end-to-end prediction accuracy, we compute the relative, reconstructed in the ambient-space, $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ per case $c$ in the testing set and time step $k$ as:
@@
For a summary of statistics, we report the mean and the 10–90% percentiles of the above errors over all cases and time steps. To track down the prediction accuracy in time, we further depict the mean and 10–90% percentiles of the error evolution in time over all cases. For the counterflow, these errors are computed separately for each group $l=1,2$.
All numerical computations were executed on a PC equipped with an 11th Gen Intel(R) Core(TM) i7-1165G7 2.80 GHz CPU and 16 GB of system memory. The source code is publicly available at github.com/patsatzisdim/NGEF_CD.